OpenStudy (anonymous):

A pair of sandals is on sale for 20% off the original price. If the original price is $16.00, what is the sale price?

OpenStudy (anonymous):

16 - (20* 16 )/ 100 = 16 - 16/5 = 16 - 3.2 = $14.80

OpenStudy (anonymous):

20 % of $16 0.2 * 16 = 3.2 and then subtract that from $16 and get sale price = $12.80

OpenStudy (anonymous):

right - i cant subtract! - it is 12.80

OpenStudy (anonymous):

thank you but i understand it more if you would set up a proportion for me ? so i could understand beacuse i am confused

OpenStudy (anonymous):

???????????

OpenStudy (anonymous):

what do you mean by proportion

OpenStudy (anonymous):